FEL Upgrade Project Weekly Brief
July 23-27, 2001

Jefferson Lab and the scientific community were saddened this week by the passing of our dear friend
and colleague, Nathan Isgur, who lost a courageous battle with cancer.  For the last decade Nathan
served as Jefferson Lab’s Chief Scientist, Head of the Nuclear Theory Group and the University
Relations Office.  A memorial service is being planned for September in Williamsburg.  Information will
be posted on the following website: http://www.jlab.org/Isgur/.

This week the FEL was used to continue Gunter Leupke’s superb series of measurements on the
nature of hydrogen defects in silicon and to continue our measurements of FEL beam quality. (See the
“Operations Section).  Next week we will take up more harmonic generation experiments with our
collaborator from Aerospace Corporation (Henry Helvajian).

We have submitted contractual documents to the DOE Site Office that will allow the funding to
commence for the JTO funded laser-materials studies as of July 31 and for USAF funded UV FEL
project for expected start-up in August.  We learned this week that the funding for the UV project has
been released to the AFRL Program Office.

WBS 3 (Beam Physics):
Driver issues for MW class performance was documented (JLAB-TN-01-034).  Octupole
requirements were developed (JLAB-TN # not yet assigned).

WBS 4 (Injector):
The gun chamber drawings have been signed off and are in the shop for fabrication. The top hat flange
assembly drawings were also signed and submitted.

Gun HVPS - Progress is being made with the design of the new Gun Tank.  It will be a cylinder 70" in
diameter and will not require any removal of the short concrete wall or pad.  All of the parts of the
HVPS returned from Glassman are being identified for their mounting in the tank.

WBS 6 (RF):
Zone 3 - The HVPS for this zone was installed and is being connected.  The circulators, waveguides,
couplers, shorts and test couplers were also installed this week.  Connection to AC power will wait
until after the summer run is over.

Zone 4 - No change.  Waiting for 7 RF Control modules, the installation of the last circulator, and the
completion of the punch list.

Quarter Klystrons - The factory tests may have to be postponed 1 week to allow CPI time to test the
first article before we witness the tests.  They do not anticipate any additional problems.

Quarter HVPS - We are ready to place the orders for the upgrade transformer, inductors, rectifier
assemblies, and SCR controller.

WBS 8 (Instrumentation):
A power supply box was built for the detector Gwyn Williams will be using for the terahertz radiation
measurements being conducted.  Some cable modifications were completed to detectors that will be
used by Steve Benson during his experiment on mirror heating/distortion.
We worked on the installation of the intercom system.  There are still some cables that need to be
pulled in labs that are currently being used, the Optics Control Room, Drive Laser Clean Room and the
Control Room.  Installation will not be completed until after the summer run.
Several hours were spent working with the new OBPM design.  There are still some design issues that
need to be addressed before this system will become fully functional.  Absolute value/Sample and Hold
boards for the system have been designed and prototype boards are on their way.  An associated 32
channel Instrument Amplifier board has also been designed and the artwork completed.  Prototypes
should be ordered next week.
A programmable delay card for the new pulse controller has been designing has been built and is in the
final stages of testing.  Several VME GPVS cards were successfully tested as well.
The task system on http://laser has been updated grouping tasks by system for ease of locating specific
tasks and historical archiving.  The website has also been updated to include a cable database.  Both of
these items will continue to be upgraded as more information becomes available.
We regretfully say goodbye to James Jones as he leaves us for the Baltimore area.  He will remain on
as a casual employee for the time being.  Good luck James.
No new drawings were provided to EECAD this week.  Those completed were:  Programmable Delay
Card Schematic, Fabrication and Artwork; Pyro-electric Detector Buffer/Amplifier Schematic,
Fabrication, Assembly and Artwork and the revision of the Beamviewer Solenoid Box Wiring

WBS 9 (Transport):
Optical Chicane Dipoles (GW)
o  The bid documents are ready to be sent out by Procurement.
Injector Dipoles (DU/DV)
o  The revisions to the drawings of the Injector Dipoles (GU & GV) are going through internal review
    at DULY and have started to be sent to us for a final check.
Arc Dipoles (GY, GX, GQ)
o  AES continued design of the Reverse Bend (GQ).  Field clamps were completed, the position
    adjusters placed and assembly refinement started.
o  In magnetic modeling, the 180 Dipole (GY) is going through final adjustment that hinged on the
    design of the GX.  The bulk field and current is being matched to the GX by adjusting the Purcell gap
    and the gap.
    Following that, AES is checking the saturation due to our reduced field path length correctors and
    will then adjust the end face position and the angle of the path length coil slot to bring the field integrals
    throughout the good field region into 180 degrees of effective length.
QX (3 inch quad)
o  We are sending steel to New England Techni-Coil to start core production.
o  Magnet Enterprises International is perfecting their winding technique on the first article coil.
QT (Trim Quad)
o  All the pieces of the 3D drawing model were modified to the narrow configuration as a precursor to
    starting on the 2D detail drawings.
o  No progress as Robin Wines was handling the Octupole engineering service task order with DULY
    Research and generating a final GW magnetic model.
o  DULY Research responded to the task order for magnetic modeling and design.  In response, we
    are refining our specs and reducing the task to magnetic analysis until we understand more about the final
    configuration. This will allow DULY to get a better picture and be more realistic with regard to the
    drafting design task.
o  Dave Douglas refined his specification for the octupole in a soon to be published tech note.  The final
    result is a further reduction in field by over half but higher quality, going from 10% to 1%.

Beam Line and Vacuum
o  Work continued on the Interface Control Document for the task definition for the Arc and Optical
    Chicane chambers.  Better definition of the optical requirements lead to less drastic requirements on the
    specialty chambers through which the optical beam traverses.
o  The bid package drawing set for the girders in the return leg are nearing completion.
o  We generated the first pass at the paperwork for the CRADA for the collaboration between
    Jefferson Lab and STI Optronics to test out their compact 180 degree bend design on the FEL
    Upgrade machine.

WBS 10 (Wiggler):
The corrector coils for the wiggler are wired up.  The wiggler is now ready for measurements.  The
measurement apparatus should be done early next week.  It must then be calibrated before
measurements can start.  That is expected to happen in the middle of August.

The dispersion section vacuum chamber is ready for welding.  The machine shop which fabricated it is
now working on the wiggler chambers.  We hope to have the dispersion section chamber ready for the
final measurements of the dispersion section.

WBS 11 (Optics):
Design progress for the optical cavity chambers continues.  We have made a decision to purchase the
UHV-compatible actuators for the mirrors, and a PR will be placed today (7/27).
We did our first measurements of absorption levels in our 6 micron HR mirrors in the mirror test stand
(MTS).  Data analysis remains to be done, but IR thermometry results are consistent with a loss less
than 500 ppm.

Provided optics hardware and support to G. Luepke (College of William and Mary).  A beautiful set of
data was obtained on H-lifetime vs. temperature from 77k to RT. This continues Prof. Luepke’s string
of publication quality measurements every time he runs.

We began checking equipment and hardware/software for measuring phase noise.  Optimized
alignment of the FEL autocorrelator, and found zero-crossing point.  Will continue testing when we
have beam upstairs next week.  Installed hardware to measure FEL laser mode and beam quality at the
FEL.  This will be removed at the end of the day.  We have begun setting up hardware for next week's
experiments (materials processing, UV generation, and micromachining).